Grade 12th passMechanics

Q .1- A projectile projected at some angle with vel. 50m/s crosses a 20m wall after 4s from the time of projection. The angle of projection of particle will be ? ( pls give full sol.) 2. a swimmer can swim in still water at 1.5m/s , the river is flowing with 2m/s from left to right. The minimum time taken by swimmer to cross the river and drift will be? ( pls full sol. )

AJAY KUMAR
12 Years agoGrade 12th pass

1 Answer

Sumit Majumdar
12 Years ago
Solution:
  1. Let the angle of projection be \Theta. Then using the expression of vertical height covered, we get:
20=\frac{50^{^{2}}sin^{2}\Theta }{2g}
On solving this we get:
sin\Theta =\frac{2}{5}
So, \Theta =sin^{-1}\left ( {\frac{2}{5}} \right)
  1. The distance between the left to right end of the river isnt given.
Thanks & Regards
Sumit Majumdar,
askIITians Faculty
Ph.D,IIT Delhi